Evaluate the limit "\displaystyle{\lim\limits_{x \to 0} \dfrac{e^{x}-x-1}{x^2}}" by using l'Hopital's Rule twice.


1
Expert's answer
2021-09-21T00:03:57-0400

"\lim\limits_{x\rightarrow0} \frac{e^x-x-1}{x^2}"

This is a "\frac{0}{0}" form, so applying L'Hopital's rule here, we get:

"=\lim\limits_{x\rightarrow0} \frac{e^x-1}{2x}"

Again, this is a "\frac{0}{0}" form, so applying L'Hopital's rule here, we get:

"=\lim\limits_{x\rightarrow0} \frac{e^x}{2}"

"=\frac{e^0}{2}\\
=\frac{1}{2}"